OpenStudy (anonymous):

what is the limit of (3+h)^-1-3-1/h as h approaches 0?

OpenStudy (anonymous):

is it maybe \[\frac{(3+h)^{-1}-\frac{1}{3}}{h}\]

OpenStudy (anonymous):

no my fault its (3+h)^-1+3^-1/h

OpenStudy (anonymous):

ok well that is fine, because it is the same thing, as \(3^{-1}=\frac{1}{3}\)

OpenStudy (anonymous):

in fact the first step is to get rid of the negative exponent, and do some algebra

OpenStudy (anonymous):



OpenStudy (anonymous):

no

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

lets go slow and also lets forget the \(h\) in the denominator for a second

OpenStudy (anonymous):

\[(3+h)^{-1}=\frac{1}{3+h}\] and \[3^{-1}=\frac{1}{3}\] so your numerator is \[\frac{1}{3+h}-\frac{1}{3}\]

OpenStudy (anonymous):

first job is to subtract (we will come back to the \(h\) in the denominator later)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

get a common denominator

OpenStudy (anonymous):

yeah and don't multuply out, leave in factored form you get \[\frac{3-(3+h)}{3(3+h)}\]

OpenStudy (anonymous):

as always, everything in the top without an \(h\) goes, leaving you with \[\frac{-h}{3(3+h)}\]

OpenStudy (anonymous):

now recall that there was an \(h\) in the bottom so we put that back

OpenStudy (anonymous):

or just cancel, either way \[\frac{-h}{3(3+h)h}=\frac{-1}{3(3+h)}\]

OpenStudy (anonymous):

let me know if and where i lost you, because there is only one more step, namely take the limit by replacing \(h\) by \(0\)

OpenStudy (anonymous):

i got it thank you... have a tough time with algebra rules that havent been learned in years

OpenStudy (anonymous):

yeah don't fret too much this is the first or second week of calculus by the third week you will say "this is the derivative of \(\frac{1}{x}\) which is \(-\frac{1}{x^2}\) if \(x=3\) so the answer has to be \(-\frac{1}{9}\)

OpenStudy (anonymous):

haha hopefully i get that good